OpenStudy (mythorius):

Click an item in the list or group of pictures at the bottom of the problem and, holding the button down, drag it into the correct position in the answer box. Release your mouse button when the item is place. If you change your mind, drag the item to the trashcan. Click the trashcan to clear all your answers. Using the technique in the model above, find the missing side in this 30°-60°-90° right triangle. Short = 5 , hypotenuse = 10 Long =

OpenStudy (mathstudent55):

You need to know the ratios of the lengths of the sides of a 30-60-90 triangle. \(1 : \sqrt 3 : 2\)

OpenStudy (mythorius):

@KamiBug

OpenStudy (mythorius):

And @mathstudent55 That's confusing, math isn't my best subject

OpenStudy (mathstudent55):

Ok. I'll explain.

OpenStudy (mythorius):

That would be nice, thank you.

OpenStudy (mathstudent55):

As you can see in your drawing, a 30-60-90 triangle is a right triangle. It has two legs, the sides that form the right angle, and a hypotenuse, the side opposite the right angle.

OpenStudy (mathstudent55):

The legs are of different lengths. The leg opposite the 30-degree angle is shorter than the leg opposite the 60-degree angle.

OpenStudy (mathstudent55):

In addition, in a right triangle, the hypotenuse is always the longest side.

OpenStudy (mathstudent55):

The three sides, in order of length, from shortest to longest are: short leg, long leg, hypotenuse

OpenStudy (mathstudent55):

Do you understand so far?

OpenStudy (mythorius):

Okay I understand the angles of triangles, I know a bit of geometry but this is tough for me.

OpenStudy (mathstudent55):

Just read the above again. So far it's pretty simple. In a 30-60-90 triangle, the shortest side is the short leg (opposite the 30-deg angle). Then the next longer side is the long leg (opposite the 60-deg angle). The longest side is the hypotenuse (opposite the right angle).

OpenStudy (mythorius):

Okay. I know that. but the confusing part is the equation part. Not the triangle.

OpenStudy (mythorius):

Is it 5 (square root) 3?

OpenStudy (mathstudent55):

Now that we know the order of the lengths of the sides, we can go one step further. The lengths of the sides of a 30-60-90 triangle are in the ratio: \(1 : \sqrt 3 : 2\) If the shortest leg measures 1, then the hypotenuse measures 2. That means the hypotenuse is twice the length of the short leg. The long leg measures \(\sqrt 3\). That means the long leg is \(\sqrt 3\) times the length of the short leg.

OpenStudy (mythorius):

Yeah so the long leg would be 15. but what would the equation be? 5x^2+10x^2?

OpenStudy (mathstudent55):

You are correct. Since the4 short leg is 5, the long leg is \(\sqrt 3\) time 5, so it is \(5 \sqrt 3\)

OpenStudy (mythorius):

Alright, that makes sense, thank you.

OpenStudy (mathstudent55):

You can also use the Pythagoras theorem. \(a^2 + b^2 = c^2\) \(5^2 + b^2 = 10^2\) \(25 + b^2 = 100\) \(b^2 = 75\) \(b = \sqrt {75} \) \(b = \sqrt {25 \times 3} \) \(b = 5\sqrt 3\)

OpenStudy (mythorius):

Can you help me with some more?

OpenStudy (mathstudent55):

Sure. Please start a new post for each new question.

OpenStudy (mythorius):

I know

OpenStudy (mathstudent55):

You're welcome.